OpenStudy (anonymous):

Describe how to transform _____ into an expression with a rational exponent.

OpenStudy (anonymous):

\[\huge \sqrt[n]{x^m}=x^{\frac{m}{n}}\]

OpenStudy (anonymous):

\[(^3 \sqrt{} x^4) ^5\]

OpenStudy (anonymous):

So do I add 5 and 4

OpenStudy (anonymous):

and then get 9 over 3?